OpenStudy (brucebaner):

Which expressions could complete this equation so that it has an infinite number of solutions? Choose exactly two answers that are correct.

OpenStudy (jdoe0001):

have you covered :system of equations" yet?

OpenStudy (brucebaner):

A \[4x-14+8x\] B\[12x-14\] C\[6x+2(3x-8)\] D\[7(x-2)+6x\]

OpenStudy (brucebaner):

Which expressions could complete this equation so that it has an infinite number of solutions? \[4(3x-5)+6\] Choose exactly two answers that are correct.

OpenStudy (jdoe0001):

have you covered :system of equations" yet?

OpenStudy (brucebaner):

yes

OpenStudy (jdoe0001):

if 2 equations "simplify" to exactly the same thing that means that they're really both exactly the same but in disguise and when the system of equations have 2 who "simplify" to the same equation, they have infinite solutions for example say 2x+3 = 5 and 4x-7 = 11 notice the 2nd one 4x -7 = 11 => 4x =4 => \(\bf x = 1\) notice the 1st one 2x +3 = 5 => 2x = 2 => \(\bf x = 1\)

OpenStudy (jdoe0001):

they look different, but they simplify to the "same", so they're really the same in disguise so..... any ideas on what of your choices is really "4(3x-5) + 6" in disguise?

OpenStudy (jdoe0001):

wel.... actaull above should have been 4x +7 = 11 anyhow 4x + 7 = 11 => 4x = 4 = > \(\bf x =1\)

OpenStudy (jdoe0001):

can you simplify "4(3x-5) + 6" ? what would it give you?